Initially the sample contained 2.0 g of the radioactive isotope. Calculate the number of radioactive nuclei in the sample after 210 seconds.

The number of nuclei in 262 g of dubnium is 6.02 x 1023

(decay constant: 0.02s-1)


1
Expert's answer
2021-02-28T07:23:17-0500

"N = N_0\times e^{-λt} \\

N_0 = 2.0 \;g \\

t = 210 \;s \\

λ = 0.02 \;s^{-1} \\

N = 2.0\times e^{-0.02 \times 210} \\

= 0.0299 \;g"

Proportion:

262 g – "6.02 \times 10^{23}"

0.0299 g – x

"x = \frac{0.0299 \times 6.02 \times 10^{23}}{262} = 6.87 \times 10^{19} \;nuclei"

Answer: "6.87 \times 10^{19} \;nuclei"